A player passes a basketball to another player who catches it at the same level from which it was thrown. The initial speed of the ball is 7.4 m/s , and it travels a distance of 4.0 m .

Part A: What was the initial direction of the ball? Express your answer using two significant figures. If there is more than one answer, separate them by a comma.

Part B: What was its time of flight?

For Part A, I got 22.9 ∘ to the horizontal, but it's marking it incorrect. I don't know what I'm doing wrong.
Answer #1


R = 4 m , u = 7.4 m/s

R = u^2sin(2theta)/g

4 = 7.4^2sin(2*theta)/9.8

theta = 23 degrees, 67 degrees

T = 2usin(theta)/g

T = 2*7.4*sin(23)/9.8

T = 0.59 s, 1.4 s